# Training and Evaluating with Human Label Variation: An Empirical Study

### v3

@@ 5.3 Discussion
ReL and SL appear to perform best as a training method, and that is somewhat surprising because they are the most straightforward methods to incorporate HLV in model training. That said, our finding that SL is a strong performer is in line with what Uma et al. (2021) discovered. The results of ReL, however, stand in contrast with prior work (Uma et al., 2021; Kurniawan et al., 2024). This difference may be due to the structured prediction tasks that they considered which have an exponentially large output space. We note that Uma et al. (2021) also found that ReL performs really well in simple, multiclass classification tasks, in line with our findings.
@@ 6 Meta-Evaluation of HLV Evaluation Metrics
In the previous section, we gained an understanding about the best training methods that perform consistently well across metrics. In this section, we tackle the question of what the best evaluation metric is for HLV data. We address this question by conducting a meta-evaluation of all the evaluation metrics in question. Intuitively, a good evaluation metric should correlate well with human preference. That is, if a model has a high performance according to a certain metric, we expect human judges to prefer that model’s predictions. This forms the basis of our meta-evaluation experiment.
To obtain a reliable estimate of human preference, we rely on subject matter experts. This is so that when they disagree we can be confident that the disagreement is a result of genuine variation rather than noise. In our case, we have access to practising lawyers who are more than qualified for our task (i.e. legal area classification). As such, we use the TAG dataset to conduct our meta-evaluation experiments despite the inability to distribute it.
@@ 6.1 Experimental Setup
For each pair of pretrained model and training method, we randomly sample an average of 28 test instances from TAG. For each instance, we (a) show the instance text, the ground truth label distribution and the predicted label distribution from one of our trained models, and then (b) ask lawyers which distribution151515Because it is multilabel task, the distribution doesn’t sum to one. is more accurate. The lawyers can choose exactly one, both, or neither of the visualisations of the distributions. On average, each instance is annotated by 3.8 lawyers. To make sure the lawyers form their own expectation on the areas of law relevant to an instance, we first show the instance text and ask them to annotate it (Figure 5). After completing this task, we then show them the label distributions from the model and ground truth (Figure 6). We conduct several pilot studies with the lawyers to refine the interface design to ensure the task is clear to our target annotators.
It is worth noting that the meta-evaluation task inherently has HLV as there is not a single, objective ground truth. Every lawyer has their own preference, and our goal is to capture the collective preference of the lawyers (i.e. the proportion of lawyers who prefer the predicted distributions). To minimise the risk of lawyers figuring out which of the two judgement distribution options is the ground truth distribution, we randomise the options and the order of the instances given to the lawyers. To filter out low quality predictions, we only show areas of law whose judgement exceeds a threshold of .161616We arrive at this threshold by choosing a value that results in similar numbers of labels whose judgement exceeds the value in the predicted and the ground truth distributions.
@@ 6.2 Results
We meta-evaluate HLV evaluation metrics by computing the correlation between their value and the fraction of times the lawyers select the predicted judgement distribution. Figure 7 shows that all metrics exhibit a linear relationship with human preference, which is positive for all metrics other than entropy correlation, suggesting that it is a poor HLV metric.
Table 5 confirms that all metrics are positively correlated with human preference except for entropy correlation. Furthermore, it shows that only soft micro F1 and PO-JSD correlate with human judgements at a level of statistical significance. This finding suggests that both metrics are good for evaluation in the HLV context.
6.3 Method ranking based on human preference
Given the human annotations above, we can compute the performance of an HLV training method based on the number of times humans select the predicted judgement distribution. Table 6 shows that the rankings of the methods have some similarities to TAG results in Figures 2, 3 and 4. For instance, ReL, JSD and SL are generally strong performers (exception: SL with LLaMA). That said, we also see some discrepancies. For example, AR and LA-min come out on top here but this is not reflected in Figures 2, 3 and 4. We contend, however, that the meta-evaluation study is ultimately based on one dataset and so we should interpret these results cautiously.
| Type | Metric | Pearson |
| Hard | Micro F1 | |
| Macro F1 | ||
| Soft | Entropy correlation | |
| PO-JSD | ∗ | |
| Soft micro F1 | ∗ | |
@@ 6.3 Method ranking based on human preference
| RoBERTa | LLaMA | ||
| Rank 1–7 | Rank 8–13 | Rank 1–7 | Rank 8–13 |
| JSD | SLMV | AR | MV |
| LA-min | SMF1 | SmF1 | SL |
| SL | ARh | LA-min | ARh |
| MV | AE | ReL | AEh |
| AR | LA-max | JSD | AE |
| ReL | AEh | SLMV | SMF1 |
| SmF1 | LA-max | ||
@@ 6.4 Discussion
Putting all the results together, PO-JSD and soft micro F1 are arguably some of the best metrics for HLV data. That said, Uma et al. (2021) argued that the value of PO-JSD is typically confined within a small range and lacks an intuitive interpretation. In contrast, soft micro F1 is more interpretable. Therefore, our general recommendation is to report both metrics, but focus on the soft micro F1 score when an accessible interpretation is important (e.g., communicating with non-technical people) or one is restricted to a single metric (e.g., in hyperparameter tuning).
@@ 7 Conclusions
We propose new evaluation metrics for evaluating model predictions with human label variation (HLV). Taking inspiration from remote sensing research, we represent human judgement distributions as degrees of membership over fuzzy sets and generalise standard metrics such as accuracy into their soft versions using fuzzy set operations. Therefore, our proposed metrics have intuitive interpretations and reduce to standard hard metrics if there is no label variation. We also generalise both our proposed metrics and some of existing ones to multilabel classification settings. We mathematically prove that our proposed soft accuracy metric is upper-bounded by an existing metric based on the Jensen-Shannon divergence.
Because the soft metrics are differentiable, we propose 3 new training methods for HLV using the metrics as the training objective. Additionally, we propose 2 new training methods that aggregate losses over annotations of the same instance. We test our proposed methods on 6 datasets spanning binary, multiclass, and multilabel classification tasks, as well as crowd and expert annotators. We evaluate across 2 language models and using a total of 6 HLV evaluation metrics including both existing and proposed ones. We compare against 9 existing HLV training methods and find that training on disaggregated annotations (ReL) or soft labels (SL) performs best across evaluation metrics, and that the former is likely to perform better than the latter when both the number of annotations per instance is large and a large language model is used. Our meta-evaluation of these metrics shows that our proposed soft metric correlates best with human preference. We recommend further research to include our proposed soft metric when reporting model performance in the HLV context.
@@
where is a function that satisfies 3 properties: commutativity, associativity, and monotonicity.171717 where The function is called a t-norm (triangular norm). A special t-norm is the function because it is also idempotent, i.e. .
@@
Taking the mean of Equation 3 over classes results in the soft macro F1 score. To obtain the micro variant, we simply modify the sums in both the numerator and the denominator of Equations 1 and 2 to also iterate over classes to get the soft micro precision and recall scores, and then take the harmonic mean of the two scores as normal.
@@
Table 7 reports the performance of MV across datasets and evaluation metrics.
@@ Acknowledgements.
This research is supported by the Australian Research Council Linkage Project LP210200917181818https://dataportal.arc.gov.au/NCGP/Web/Grant/Grant/LP210200917 and funded by the Australian Government. This research is done in collaboration with Justice Connect, an Australian public benevolent institution.191919As defined by the Australian government: https://www.acnc.gov.au/charity/charities/4a24f21a-38af-e811-a95e-000d3ad24c60/profile We thank Kate Fazio, Tom O’Doherty, and Rose Hyland from Justice Connect for their support throughout the project. We thank David McNamara, Yashik Anand, and Mark Pendergast also from Justice Connect for the development of the meta-evaluation annotation interface. This research is supported by The University of Melbourne’s Research Computing Services and the Petascale Campus Initiative.References

### v4

@@ 5.3 Discussion
ReL and SL appear to perform best as a training method, and that is somewhat surprising because they are the most straightforward methods to incorporate HLV in model training. That said, our finding that SL is a strong performer is in line with what Uma et al. (2021) discovered. The results of ReL, however, stand in contrast with prior work (Uma et al., 2021; Kurniawan et al., 2024). This difference may be due to the structured prediction tasks that the prior work considered which have an exponentially large output space. We note that Uma et al. (2021) also found that ReL performs really well in simple, multiclass classification tasks, in line with our findings.
@@ 6 Meta-Evaluation of HLV Evaluation Metrics
In the previous section, we saw the best training methods that perform consistently well across metrics. In this section, we address the next question: what is the best evaluation metric for HLV data? We answer this by conducting a meta-evaluation on the evaluation metrics. We focus on the TAG dataset specifically for this meta-evaluation experiment as we have access to practicing lawyers as annotators who can provide high quality annotations. It does mean we are unable to distribute the dataset for ethical reasons, but we believe the insights derived from the meta-evaluation is worth the trade-off.
A good evaluation metric should produce a ranking of training methods (henceforth “metric ranking”) that correlates to the ranking produced by human judgements (henceforth “human ranking”). To create the latter (human ranking), we frame it as a pairwise comparison task where we pit two methods against each other and ask lawyers to select one of them. Given a large number of human-judged pairwise comparisons, we then use an algorithm to create a ranking for the training methods. Intuitively, the algorithm will rank a method that “wins” consistently higher than another method that “loses” most of the time. Once we have the human ranking, we can then compute its correlation with each metric ranking, and the best metrics are the ones with the highest correlations.
@@ 6.1 Experimental Setup
To create the ranking of training methods produced by human judgements, given a language model (RoBERTa or LLaMA) we randomly sample a pair of methods (e.g. ReL and SL) and present the instance text (i.e. a legal problem) and two results (i.e. distribution of the areas of law)191919Note that because it is multilabel task, the distribution doesn’t sum to one. We only show areas of law whose probability exceeds a threshold of . We choose this threshold because it results in similar numbers of areas of law whose probability exceeds the value across all methods. produced by the two methods, and then ask the lawyers which result is more accurate. The lawyers can choose either one of them, both, or neither of them. See Appendix Figure˜7 for an illustration of the annotation task. In total, we collect 1.2K pairwise judgements for both RoBERTa and LLaMA. On average, each instance is annotated by 2.3 lawyers and each pair of methods has 8 judgements.
To make sure the lawyers form their own expectation on the areas of law relevant to an instance before making this pairwise selection, we first show the instance text and ask them to select the most relevant areas of law (Figure˜6 in the Appendix). After completing this task, we then ask them to make the pairwise selection. We conduct several pilot studies with the lawyers to refine the interface design and ensure the task is clear to them.
At the end of the annotation task, we have a collection of human-judged pairwise comparisons for the training methods. We next use the rank centrality algorithm (Negahban, Oh, and Shah, 2012) to compute the “score” for each method to ultimately produce a ranking of the methods. This algorithm casts the problem as a random walk on a weighted directed graph where each node corresponds to a method and the weight of the edge from method to method is the probability that method wins over model .202020We determine that method wins over method if the annotator select method but not method . The score of a method is then its stationary probability for this random walk. Intuitively, the score is high if it wins against other high-scoring methods or against many low-scoring methods. To assess the evaluation metrics and understand which metrics are the best, for each metric we compute the Pearson correlation between: (1) method performance as given by the evaluation metric; and (2) the method scores produced by human judgements.
@@ 6.2 Results
Figure˜5 shows that not all metrics exhibit a strong linear relationship with human judgements. For example, both hard F1 scores and entropy correlation show almost no relationship for LLaMA. All metrics except entropy correlation show a strong linear relationship for RoBERTa. This suggests that entropy correlation is a poor HLV metric.
Table˜5 confirms that all metrics are positively correlated with human judgements but with different magnitudes. As before, both hard F1 scores and entropy correlation are very weakly correlated with human judgements for LLaMA, and entropy correlation is the weakest for RoBERTa. Furthermore, it is the only metric whose correlation is not statistically significant. Across both pretrained models, both PO-JSD and soft micro F1 consistently have the strongest correlation, but each is best for different models: PO-JSD is best for LLaMA while soft micro F1 score is best for RoBERTa. This finding suggests that both metrics are good for evaluation in the HLV context.
6.3 Method ranking based on human judgements
Given the scores computed from human judgements, we can rank the methods based on their scores (human ranking). Table˜6 shows that the human rankings have some similarities to TAG results in Figure˜4. For instance, ReL, JSD and SL are generally strong performers (exception: ReL with LLaMA). That said, we also see some discrepancies. For example, SLMV is highly ranked here but not reflected in Figure˜4. We contend, however, that the meta-evaluation study is ultimately based on one dataset and so we should interpret these results cautiously.
| Model | Type | Metric | Pearson | -value |
| RoBERTa | Hard | Micro F1 | ∗ | |
| Macro F1 | ∗ | |||
| Soft | Entropy correlation | |||
| PO-JSD | ∗ | |||
| Soft micro F1 | ∗ | |||
| Soft macro F1 | ∗ | |||
| LLaMA | Hard | Micro F1 | ||
| Macro F1 | ||||
| Soft | Entropy correlation | |||
| PO-JSD | ||||
| Soft micro F1 | ||||
@@ 6.3 Method ranking based on human judgements
| RoBERTa | LLaMA | ||||||
| Rank 1–7 | Rank 8–13 | Rank 1–7 | Rank 8–13 | ||||
| Method | Score | Method | Score | Method | Score | Method | Score |
| SMF1 | MV | SL | ReL | ||||
| JSD | SmF1 | SLMV | LA-max | ||||
| ARh | LA-max | AE | ARh | ||||
| SLMV | AR | MV | LA-min | ||||
| SL | AE | AR | AEh | ||||
| ReL | AEh | JSD | SMF1 | ||||
| LA-min | SmF1 | ||||||
@@ 6.4 Discussion
Putting all the results together, PO-JSD and soft micro F1 are arguably some of the best metrics for HLV data. That said, Uma et al. (2021) argued that the value of PO-JSD is typically confined within a small range and lacks an intuitive interpretation. We believe that this is its biggest weakness, which is shared by other information-theoretic measures such as cross-entropy. In contrast, soft micro F1 is more interpretable as it is analogous to the standard F1 score. Therefore, our general recommendation is to report both metrics, but focus on the soft micro F1 score when an accessible interpretation is important (e.g., communicating with non-technical people) or one is restricted to a single metric (e.g., in hyperparameter tuning).
@@ 7 Conclusions
We propose new evaluation metrics for evaluating model predictions with human label variation (HLV). Taking inspiration from remote sensing research, we represent human judgement distributions as degrees of membership over fuzzy sets and generalise standard metrics such as accuracy into their soft versions using fuzzy set operations. Therefore, our proposed metrics have intuitive interpretations and reduce to standard hard metrics if there is no label variation. While our analysis shows that our proposed soft accuracy metric is strongly correlated with an existing metric based on Jensen-Shannon divergence, we mathematically prove that the former is upper-bounded by the latter. We further show that the JSD-based metric can produce a misleadingly high score as a result.
Because the soft metrics are differentiable, we propose 3 new training methods for HLV using the metrics as the training objective. Additionally, we propose 2 new training methods that aggregate losses over annotations of the same instance, bringing the total of our proposed training methods to 5 methods. We test our proposed methods on 6 datasets spanning binary, multiclass, and multilabel classification tasks, as well as crowd and expert annotators. We evaluate against 9 existing HLV training methods across 2 pretrained models and use a total of 6 HLV evaluation metrics including both existing and proposed ones to find the best methods for HLV data. Then, we perform an empirical meta-evaluation of the evaluation metrics to understand which metrics are best for HLV data.
We find that simple methods such as training on disaggregated annotations (ReL) or soft labels (SL) perform best in most cases. They often outperform not only training on the majority-voted labels but also more complex HLV training methods including our proposed training methods with the differentiable metrics. Our meta-evaluation shows that our proposed soft micro F1 score is one of the best metrics for HLV data. This metric reduces to our proposed soft accuracy in single-label classification. Given its intuitive interpretation and positive meta-evaluation result, we recommend further research to include soft micro F1 when reporting model performance in the HLV context.
@@
where is a function that satisfies 3 properties: commutativity, associativity, and monotonicity.212121 where The function is called a t-norm (triangular norm). A special t-norm is the function because it is also idempotent, i.e. .
@@
Taking the mean of Equation˜3 over classes results in the soft macro F1 score. To obtain the micro variant, we simply modify the sums in both the numerator and the denominator of Equations˜1 and 2 to also iterate over classes to get the soft micro precision and recall scores, and then take the harmonic mean of the two scores as normal.
@@
Table˜7 reports the performance of MV across datasets and evaluation metrics.
@@
Annotation Interface Examples
Figures˜6 and 7 show the interface of the area of law and the preference annotation tasks respectively.
@@ Acknowledgements.
This research is supported by the Australian Research Council Linkage Project LP210200917222222https://dataportal.arc.gov.au/NCGP/Web/Grant/Grant/LP210200917 and funded by the Australian Government. This research is done in collaboration with Justice Connect, an Australian public benevolent institution.232323As defined by the Australian government: https://www.acnc.gov.au/charity/charities/4a24f21a-38af-e811-a95e-000d3ad24c60/profile We thank Kate Fazio, Tom O’Doherty, and Rose Hyland from Justice Connect for their support throughout the project. We thank David McNamara, Yashik Anand, and Mark Pendergast also from Justice Connect for the development of the meta-evaluation annotation interface. This research is supported by The University of Melbourne’s Research Computing Services and the Petascale Campus Initiative.References
@@
- Negahban, Oh, and Shah (2012) Negahban, Sahand, Sewoong Oh, and Devavrat Shah. 2012. Iterative ranking from pair-wise comparisons. In Advances in Neural Information Processing Systems, volume 25.
@@
- Rizzi et al. (2024) Rizzi, Giulia, Elisa Leonardelli, Massimo Poesio, Alexandra Uma, Maja Pavlovic, Silviu Paun, Paolo Rosso, and Elisabetta Fersini. 2024. Soft metrics for evaluation with disagreements: An assessment. In Proceedings of the 3rd Workshop on Perspectivist Approaches to NLP (NLPerspectives) @ LREC-COLING 2024, pages 84–94.